1.A Particle moves in a curve y=A LOG(secx/a) such that tangent to the curve rotates uniformly > prove that the resultant acceleration of the particle varies as the square of the radius of curvatur

Grade:12th pass

1 Answers

Arun
4 years ago
Y = ln(sec(x)/a),Putting a =1,Y=ln(sec(x)),Slope of this curve,Tan(θ)=dy/dx=tan(x)........equation(1),θ=x,Angular velocity of tangent i.e. ω=dθ/dt=dx/dt=2,Hence dx/dt=2,By equation (1),dy/dt =tan(x) dx/dt,Hence velocity along y direction=2tan(x),So acceleration along y,= 2 sec(x) sec(x),At x=π/4, acceleration=2 √2 √2 = 4